# Positive microlocal holonomies are globally regular

### v1

@@ 1.2. Main result
Acknowledgements: R. Casals thanks the hospitality of the Institute of Advanced Study at Princeton, where a significant part of this article was written. R. Casals is supported by the NSF CAREER DMS-1942363, a Sloan Research Fellowship of the Alfred P. Sloan Foundation and a UC Davis College of L&S Dean’s Fellowship. W. Li is partially supported by the AMS-Simons Travel Grant. We thank M. Christ, C. Kuo and H. Williams for helpful discussions on the $\operatorname{HO}$ map and higher category theory. $\Box$
[…]
@@ 2.1. Dg-categories of sheaves with Legendrian singular support
The dg-categories $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$ and $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ are Legendrian isotopy invariants of $\Lambda$, cf. [GKS12, Section 3]. Their information only allows for the definition of microlocal monodromies on a given exact Lagrangian filling $L$ of $\Lambda$; microlocal monodromies are defined using the fully faithful embedding $\mathrm{Loc}(L)\hookrightarrow\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$, cf. Subsection 3.4 below or [JT24]. These monodromies are only rational functions on the moduli of pseudo-perfect objects (intuitively, on the moduli of Lagrangian fillings of $\Lambda)$, and are typically not regular. In particular, it is challenging to extract geometric information from them, cf.  e.g. [CW24, Sections 2.8 & 4.4] and [CGG+22, Section 8] for further discussions.
Microlocal merodromies, allowing for more general parallel transports along relative cycles with decorated ends, have proven more useful, see e.g. [CW24, Section 4], [CGG+22, Section 5] and [CL23, Section 6]. Therefore, we now introduce in Definition 2.1, and see also Definition 2.2, an enhancement of $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$ or $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ which allows for microlocal merodromies to be defined and used. [^2] The results and arguments in this section hold for the dg-categories $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$ and $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$. In fact, by [GPS24, Cor. 4.22], the inclusion $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}\hookrightarrow\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$ has a left adjoint that preserves compact objects and thus the derived moduli stack of pseudo-perfect objects of $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ is an open and closed substack of the corresponding moduli for $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$. We state the results for $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ only since it better fits into the comparison with the Kashiwara-Schapira stacks for $\mathbb{L}$-compressing systems in Section 3.
@@
where ${\Lambda}\times\mathbb{R}_{+}\subset T^{*}\mathbb{R}^{2}$ is the conormal cone associated to $\Lambda\subset T^{*}_{\infty}\mathbb{R}^{2}$ and $m_{{\Lambda}\times\mathbb{R}_{+}}$ is the microstalk functor of ${\Lambda}\times\mathbb{R}_{+}$, e.g.  as defined in [Nad16, Definition 3.8]. See also [KS90, Section 6] for details on the microlocal theory of sheaves. Equivalently, $m_{\Lambda,\mathfrak{t}}$ assigns to an object $\mathscr{F}$ the stalks of the local system $m_{\Lambda}(\mathscr{F})\in\mathrm{Loc}(\Lambda)$ at the points in $\mathfrak{t}$, where $m_{\Lambda}:\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}\longrightarrow\mathrm{Loc}(\Lambda)$ is the microlocalization functor, cf. Definition 3.3 below or [CL23, Appendix A] and references therein. [^3] Note that here we have implicitly chosen an isomorphism $\mu\mbox{sh}_{\Lambda}(\Lambda)\cong\mathrm{Loc}(\Lambda)$ between the global sections of the Kashiwara-Schapira stack $\mu\mbox{sh}$ and the dg-derived category of local systems $\mathrm{Loc}(\Lambda)$ on $\Lambda$. This is possible in the case where $\Lambda=S^{1}\sqcup\ldots\sqcup S^{1}$ by [Gui23, Chapter 10]. This is also discussed later in Section 3. In this latter perspective, we are implicitly using an equivalence
@@
which is independent of $\Lambda$ and $T$. The microstalk functor $m_{\Lambda,\mathfrak{t}}$ admits a left adjoint $m^{\ell}_{\Lambda,\mathfrak{t}}$ because it preserves products, see e.g. [Nad16, Section 3.6] or [GPS24, Lemma 4.13]. Similarly, $\Delta$ admits a left adjoint $\Delta^{\ell}$ given by the coproduct. Both functors $m_{\Lambda,\mathfrak{t}}$ and $\Delta$ preserve coproducts, and thus $m^{\ell}_{\Lambda,\mathfrak{t}}$ and $\Delta^{\ell}$ preserve compact objects.
Consider the full dg-subcategories $\operatorname{Sh}^{c}_{\Lambda}(\mathbb{R}^{2})_{0}\subset\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ and $\operatorname{\mbox{Mod}}(k)^{c}\subset\operatorname{\mbox{Mod}}(k)$ of compact objects. The category $\operatorname{\mbox{Mod}}(k)^{c}$ is the dg-subcategory of perfect complexes, i.e. $\operatorname{Perf}(k)$. These are both smooth dg-categories, cf. [GPS24, Corollary 4.26]. In fact, $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})^{c}_{0}$ is even of finite type by the results from [Nad17, Sta18] (or by virtue of being equivalent to a wrapped Fukaya category, as proven in [GPS24]). Note that $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})^{c}_{0}$ is a small dg-category, as it is the subcategory of compact objects of the compactly generated category $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$.
@@
${\displaystyle\prod_{i=1}^{|\mathfrak{t}|}\operatorname{\mbox{Mod}}(k)}$${\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}}$${\operatorname{\mbox{Mod}}(k)}$$\scriptstyle{m_{\Lambda,\mathfrak{t}}^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
We refer to $\operatorname{Sh}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ as the category of compactly supported sheaves with singular support in $(\Lambda,\mathfrak{t})$. $\Box$
[…]
@@ 2.2. Derived stacks of sheaves with Legendrian singular support
$\mathfrak{M}_{\mathcal{C}}:\mbox{sCAlg}_{k}\to\mbox{sSet},\quad\mathfrak{M}_{\mathcal{C}}(A)=\mbox{Map}_{\operatorname{dg-cat}_{k}}(\mathcal{C}^{op},\operatorname{Perf}(A)).$
[…]
@@
[…]
@@
(2.3) ${\mathfrak{M}\left(\displaystyle\prod_{i=1}^{|\mathfrak{t}|}\operatorname{\mbox{Mod}}(k)^{c}\right)}$${\mathfrak{M}(\operatorname{Sh}^{c}_{\Lambda}(\mathbb{R}^{2})_{0})}$${\mathfrak{M}(\operatorname{\mbox{Mod}}(k)^{c})}$${{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathfrak{M}(\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0})}.}$$\scriptstyle{\mathfrak{M}(m_{\Lambda,\mathfrak{t}}^{\ell})}$$\scriptstyle{\mathfrak{M}(\Delta^{\ell})}$
@@
$\mathfrak{M}(\operatorname{Sh}^{c}_{\Lambda}(\mathbb{R}^{2})_{0})\longrightarrow\mathfrak{M}\left(\prod_{i=1}^{|\mathfrak{t}|}\operatorname{\mbox{Mod}}(k)^{c}\right)$
@@
$\mathfrak{M}(\operatorname{\mbox{Mod}}(k)^{c})\longrightarrow\mathfrak{M}\left(\prod_{i=1}^{|\mathfrak{t}|}\operatorname{\mbox{Mod}}(k)^{c}\right)$
@@
Note that $\mathfrak{M}({\Lambda,\mathfrak{t}})$ depends on the choice of $\mathfrak{t}$, e.g. the dimension of its 00-truncation $t_{0}\mathfrak{M}({\Lambda,\mathfrak{t}})$ depends on the cardinality of $\mathfrak{t}$. That said, there is a natural choice for the cardinality of $\mathfrak{t}$ given $\Lambda$: namely, we can always choose exactly one basepoint per component, i.e. $|\mathfrak{t}|=|\pi_{0}(\Lambda)|$. For each fixed microlocal rank $n\in\mathbb{N}$, the corresponding connected component of $\mathfrak{M}_{n}({\Lambda})\subset\mathfrak{M}({\Lambda})$ is a (stack) quotient of a component $\mathfrak{M}_{n}({\Lambda,\mathfrak{t}})\subset\mathfrak{M}({\Lambda,\mathfrak{t}})$ under an algebraic action. See Lemma 2.6 for details and note that this action is typically not free.

### v2

@@ 1.4. Outline of the argument
Acknowledgements: We are grateful to the referee, whose edits and suggestions have improved the article. R. Casals thanks the hospitality of the Institute of Advanced Study at Princeton, where a significant part of this article was written. R. Casals is supported by the National Science Foundation under grants DMS-2505760 and DMS-1942363, a Sloan Research Fellowship of the Alfred P. Sloan Foundation and a UC Davis College of L&S Dean’s Fellowship. W. Li is partially supported by the AMS-Simons Travel Grant. $\Box$
Notation: Given $n\in\mathbb{N}$, we denote $1,n]:=\{1,\ldots,n\}$ and $k$ denotes a commutative unital ring. Throughout the article, categories will be dg-derived categories unless otherwise specified, i.e. we work with categories enriched over chain complexes and localized at acyclic complexes. For instance, $\operatorname{\mbox{Mod}}(k)$ denotes the dg-derived category of chain complexes over $k$, and derived global sections and derived tensor product are still denoted by $\Gamma$ and $\otimes$. Similarly, $\mathrm{Loc}(\Lambda)$ denotes the dg-derived category of local systems on $\Lambda$, cf. [CL23, Appendix A] and references therein. We write $\mbox{dg-cat}_{k}$ for the $\infty$-category of small dg-categories over $k$, and $\mbox{dg-Cat}_{k}$ for the $\infty$-category of well-generated dg-categories over $k$, obtained from the corresponding model categories [TAB05, TAB09]. In terms of derived algebraic geometry, we work within the framework of $D^{-}$-stacks, cf.  [TV05, TV08], and specifically [TV07] for the derived moduli stack of pseudo-perfect objects. $\Box$
@@ 2.1. Dg-categories of sheaves with Legendrian singular support
The dg-categories $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$ and $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ are Legendrian isotopy invariants of $\Lambda$, cf. [GKS12, Section 3]. Their information only allows for the definition of microlocal monodromies on a given exact Lagrangian filling $L$ of $\Lambda$; microlocal monodromies are defined using the fully faithful embedding $\mathrm{Loc}(L)\hookrightarrow\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$, cf. Subsection 3.4 below or [JT24]. These monodromies are only rational functions on the moduli of pseudo-perfect objects (intuitively, on the moduli of Lagrangian fillings of $\Lambda)$, and are typically not regular. In particular, it is challenging to extract geometric information from them, cf.  e.g. [CW24, Sections 2.8 & 4.4] and [CGG+25, Section 8] for further discussions.
Microlocal merodromies, allowing for more general parallel transports along relative cycles with decorated ends, have proven more useful, see e.g. [CW24, Section 4], [CGG+25, Section 5] and [CL23, Section 6]. Therefore, we now introduce in Definition 2.1, and see also Definition 2.2, an enhancement of $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$ or $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ which allows for microlocal merodromies to be defined and used. [^2] The results and arguments in this section hold for the dg-categories $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$ and $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$. In fact, by [GPS24, Cor. 4.22], the inclusion $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}\hookrightarrow\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$ has a left adjoint that preserves compact objects and thus the derived moduli stack of pseudo-perfect objects of $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ is an open and closed substack of the corresponding moduli for $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$. We state the results for $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ only since it better fits into the comparison with the Kashiwara-Schapira stacks for $\mathbb{L}$-compressing systems in Section 3.
@@
where ${\Lambda}\times\mathbb{R}_{+}\subset T^{*}\mathbb{R}^{2}$ is the conormal cone associated to $\Lambda\subset T^{*}_{\infty}\mathbb{R}^{2}$ and $m_{{\Lambda}\times\mathbb{R}_{+}}$ is the microstalk functor of $\Lambda$$\times\mathbb{R}_{+}$, e.g.  as defined in [NAD16, Definition 3.8]. See also [KS90, Section 6] for details on the microlocal theory of sheaves. Equivalently, $m_{\Lambda,\mathfrak{t}}$ assigns to an object $\mathscr{F}$ the stalks of the local system $m_{\Lambda}(\mathscr{F})\in\mathrm{Loc}(\Lambda)$ at the points in $\mathfrak{t}$, where $m_{\Lambda}:\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}\longrightarrow\mathrm{Loc}(\Lambda)$ is the microlocalization functor, cf. Definition 3.3 below or [CL23, Appendix A] and references therein. [^3] Note that here we have implicitly chosen a quasi-equivalence $\mu\mbox{sh}_{\Lambda}(\Lambda)\cong\mathrm{Loc}(\Lambda)$ between the global sections of the Kashiwara-Schapira stack $\mu\mbox{sh}$ and the dg-derived category of local systems $\mathrm{Loc}(\Lambda)$ on $\Lambda$. This is possible in the case where $\Lambda=S^{1}\sqcup\ldots\sqcup S^{1}$ by [GUI23, Chapter 10]. This is also discussed later in Section 3. In this latter perspective, we are implicitly using an equivalence
@@
which is independent of $\Lambda$ and $T$. The microstalk functor $m_{\Lambda,\mathfrak{t}}$ admits a left adjoint $m^{\ell}_{\Lambda,\mathfrak{t}}$ because it preserves products, see e.g. [NAD16, Section 3.6] or [GPS24, Lemma 4.13]. Similarly, $\Delta$ admits a left adjoint $\Delta^{\ell}$ given by the coproduct. Both functors $m_{\Lambda,\mathfrak{t}}$ and $\Delta$ preserve coproducts, and thus $m^{\ell}_{\Lambda,\mathfrak{t}}$ and $\Delta^{\ell}$ preserve compact objects see e.g. [GPS24, Theorem 4.14].
Consider the full dg-subcategories $\operatorname{Sh}^{c}_{\Lambda}(\mathbb{R}^{2})_{0}\subset\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ and $\operatorname{\mbox{Mod}}(k)^{c}\subset\operatorname{\mbox{Mod}}(k)$ of compact objects. The category $\operatorname{\mbox{Mod}}(k)^{c}$ is the dg-subcategory of perfect complexes, i.e. $\operatorname{Perf}(k)$. These are both smooth dg-categories, cf. [GPS24, Corollary 4.26]. In fact, $\operatorname{Sh}_{\Lambda}^{c}(\mathbb{R}^{2})_{0}$ is even of finite type by the results from [NAD17, STA18] (or by virtue of being equivalent to a wrapped Fukaya category, as proven in [GPS24]). Note that $\operatorname{Sh}_{\Lambda}^{c}(\mathbb{R}^{2})_{0}$ is a small dg-category, as it is the subcategory of compact objects of the compactly generated category $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$.
@@
$\prod_{i=1}^{|\mathfrak{t}|}\operatorname{\mbox{Mod}}(k)$$\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$$\operatorname{\mbox{Mod}}(k)$$\scriptstyle{m_{\Lambda,\mathfrak{t}}^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
We refer to $\operatorname{Sh}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ as the category of compactly supported sheaves with singular support in $(\Lambda,\mathfrak{t})$.
$\Box$
Here the homotopy colimit is taken in the $\infty$-category $\mbox{dg-Cat}_{k}$ of well generated dg-categories over $k$, cf. [TAB09]. Intuitively, $\operatorname{Sh}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ captures compactly supported sheaves with singular support on $\Lambda$ with the additional data of their microstalks at the basepoints of $\mathfrak{t}$ and a common identification of these microstalks.
@@ 2.2. Derived stacks of sheaves with Legendrian singular support
$\mathfrak{M}_{\mathcal{C}}:\mbox{sCAlg}_{k}\longrightarrow\mbox{sSet},\quad\mathfrak{M}_{\mathcal{C}}(A)=\mbox{Map}_{\operatorname{dg-cat}_{k}}(\mathcal{C}^{op},\operatorname{Perf}(A)).$
Here $\mbox{sCAlg}_{k}$ is the category of simplicial commutative $k$-algebras (denoted $sk\mbox{-CAlg}$ in [TV07, Section 2.3]), sSet the category of simplicial sets, $\operatorname{Perf}(A)$ is the dg-category of perfect A-modules (denoted $\hat{A}_{pe}$ in [TV07, Section 2.4]), and $\mbox{Map}_{\operatorname{dg-cat}_{k}}$ denotes the mapping space of a model structure for the category of small dg-categories. See [TV07, Section 3] for the necessary details. [^4] Note that, in zero characterstic, sCAlg is equivalent to $\mbox{cdga}_{\leq 0}$, via the appropriate version of the Dold-Kan correspondence. Therefore, in zero characteristic, the inputs for (functor of points of) the derived stack $\mathfrak{M}_{\mathcal{C}}$ can be taken to be non-positively graded commutative dg-algebras. The assignment $\mathcal{C}\mapsto\mathfrak{M}_{\mathcal{C}}$ defines a functor
@@
between the opposite of the homotopy category $\mbox{Ho}(\mbox{dg-cat}_{k})$ of dg-categories and the category $D^{-}\mbox{St}(k)$ of $D^{-}$-stacks, the functor being enriched over the homotopy category of sSet. In particular, given a dg-functor $f:\mathcal{C}\longrightarrow\mathcal{D}$, there is a map $\mathfrak{M}(f):\mathfrak{M}(\mathcal{D})\longrightarrow\mathfrak{M}(\mathcal{C})$, which sends a pseudo-perfect object $\mathcal{D}^{op}\longrightarrow\operatorname{Perf}(k)$ to its pull-back to $\mathcal{C}^{op}\longrightarrow\operatorname{Perf}(k)$.
By [TV07, Prop. 3.4], the functor $\mathfrak{M}$ admits a left adjoint and therefore $\mathfrak{M}$ preserves homotopy limits. In particular, it sends a homotopy pullback in $\mbox{Ho}(\mbox{dg-cat}_{k})^{op}$ to a homotopy pullback in $D^{-}\mbox{St}(k)$. Since homotopy pullbacks in $\mbox{Ho}(\mbox{dg-cat}_{k})^{op}$ are homotopy pushouts in $\mbox{Ho}(\mbox{dg-cat}_{k})$, the functor $\mathfrak{M}$ applied to the homotopy pushout
(2.2) $\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k)$$\operatorname{Sh}^{c}_{\Lambda}(\mathbb{R}^{2})_{0}$$\operatorname{Perf}(k)$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$$\scriptstyle{m_{\Lambda,\mathfrak{t}}^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
@@
(2.3) $\mathfrak{M}\left(\displaystyle\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k)\right)$$\mathfrak{M}(\operatorname{Sh}^{c}_{\Lambda}(\mathbb{R}^{2})_{0})$$\mathfrak{M}(\operatorname{Perf}(k))$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathfrak{M}(\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0})$$\scriptstyle{\mathfrak{M}(m_{\Lambda,\mathfrak{t}}^{\ell})}$$\scriptstyle{\mathfrak{M}(\Delta^{\ell})}$
@@
$\mathfrak{M}(\operatorname{Sh}^{c}_{\Lambda}(\mathbb{R}^{2})_{0})\longrightarrow\mathfrak{M}\left(\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k)\right)$
@@
$\mathfrak{M}(\operatorname{Perf}(k))\longrightarrow\mathfrak{M}\left(\prod_{i=1}^{|\mathfrak{t}|}\operatorname{Perf}(k)\right)$
@@
Note that $\mathfrak{M}({\Lambda,\mathfrak{t}})$ depends on the choice of $\mathfrak{t}$, e.g. the dimension of its 0-truncation $t_{0}\mathfrak{M}({\Lambda,\mathfrak{t}})$ depends on the cardinality of $\mathfrak{t}$. That said, there is a natural choice for the cardinality of $\mathfrak{t}$ given $\Lambda$: namely, we can always choose exactly one basepoint per component, i.e. $|\mathfrak{t}|=|\pi_{0}(\Lambda)|$. For each fixed microlocal rank $n\in\mathbb{N}$, the corresponding connected component of $\mathfrak{M}_{n}({\Lambda})\subset\mathfrak{M}({\Lambda})$ is a (stack) quotient of a component $\mathfrak{M}_{n}({\Lambda,\mathfrak{t}})\subset\mathfrak{M}({\Lambda,\mathfrak{t}})$ under an algebraic action. See Lemma 2.3 for details and note that this action is typically not free.